# Prioritized-MVBA: A New Approach to Design an Optimal Asynchronous Byzantine Agreement Protocol

### v1

@@ 2 $\Sigma\leftarrow\{\}$
11upon receiving $\langle recommendation,ID,proposal,\sigma\rangle$ from party $p_{j}$ for the first time do
12       if  $ID.instance=instance$wedge $Verify_{id}\langle proposal,\sigma\rangle=true$  then
13             $\Sigma\leftarrow{\langle proposal,\sigma\rangle\cup\Sigma}$
14
15
16
Algorithm 2 Recommend Protocol for all parties
Detailed pMVBA Protocol We now present the full details of the pMVBA protocol, inspired by the Cachin-MVBA protocol [13] but with fewer broadcasts. The protocol replaces the $O(n)$ commit broadcast with one recommend message and runs an expected constant number of ABBA instances to reach an agreement on a party’s proposal.
The formal description of the pMVBA protocol is shown in Algorithm 3. It consists of the following steps:
  • Party/Committee Selection: When a new instance starts, a party invokes the CS protocol to select parties. (See Algorithm 1 lines 1-10 and Algorithm 3 line 10)
  • pVCBC Broadcast: A selected party broadcasts its proposal using the pVCBC protocol and generates verifiable proofs of their broadcasts. (See Algorithm 3 line 13 and Algorithm 4 lines 1-16.)
  • Propose and Recommend: Upon completing the pVCBC protocol, a party broadcasts (as a propose) its verifiable proof. When a party receives a verifiable proof, it broadcasts the proof as a recommendation and a party waits for $n-f$ recommendation messages. (See Algorithm 3 lines 14-22 and the Algorithm 2)
  • Permutation: When a party receives $n-f$ recommendations, it generates a random order of the selected parties using the cryptographic coin-tossing scheme. (See Algorithm 3 line 22 and Algorithm 4 lines 19-29 )
  • Sequential-ABBA: For the agreement phase, a party runs the ABBA protocol for the selected parties in the determined order, stopping once an agreement is reached. (See Algorithm 3 lines 32-58 and Algorithm 5 lines 1-28).
@@ 13             $\Sigma\leftarrow{\langle proposal,\sigma\rangle\cup\Sigma}$
2 Local variables initialization:
@@ 3      $instance\leftarrow 1$
4      $PParties\leftarrow\bot$
5      $\sigma\leftarrow\bot$
@@ 6      $proposal\leftarrow\bot$
7      $W\leftarrow\bot$
8      $\Sigma\leftarrow\bot$
@@ 12           $ID\leftarrow{\langle instance,id\rangle}$
13           $PParties\leftarrow CS\langle ID,instance\rangle$
@@
15          if $p_{i}\in PParties$ then
16                $proposal\leftarrow request$
17                $\sigma\leftarrow pVCBC\langle ID,requests\rangle$
[…]
@@ 17                $\sigma\leftarrow pVCBC\langle ID,requests\rangle$
20                wait for a propose or a recommend message.
21               upon receiving $\langle\textit{propose},ID,\sigma,proposal\rangle$ for the first time do
22                     $W\leftarrow SCR\langle ID,proposal,\sigma\rangle$
23
24
25                    upon receiving $\langle\textit{recommendation},ID,proposal,\sigma\rangle$ for the first time do
26                          if no $recommendation$ is sent yet  then
27                               $W\leftarrow SCR\langle ID,proposal,\sigma\rangle$
28
29
30
@@ 31                              Permutation:
32                               $\Sigma\leftarrow Permutation\langle ID,PParties\rangle$
33
@@ 34                              Sequential-ABBA:
35                               $proposal\leftarrow Sequential-ABBA\langle\Sigma,W\rangle$
@@ 37                               $instance\leftarrow instance+1$
38
@@ 5 Evaluation and Analysis of the Proposed Protocol
•  What are the formal proofs supporting the correctness of the Prioritized-MVBA protocol? (§$\S$ 5.1)
•  How does the Prioritized-MVBA protocol perform in terms of message complexity, communication complexity, and running time? (§$\S$ 5.2)
•  What are the observed behaviors and outcomes of the protocol under different network conditions and adversarial strategies? (§$\S$ 5.3)
•  How does the performance of the Prioritized-MVBA protocol compare to other MVBA protocols in terms of key metrics? (§$\S$ 5.4)
@@ 5.1 Proof of the Prioritized-MVBA Protocol
The correctness of the proposed protocol is critical to ensure that it adheres to the standard MVBA outcomes. Our goal is to prove that reducing the number of broadcasts does not compromise the protocol’s reliability. The output of a party’s request depends on the output of an ABBA instance within our protocol’s agreement loop, typically resulting in a value of 1. For the ABBA protocol to output 1, at least one honest party must input 1. We substantiate our protocol’s integrity by demonstrating that at least $12f+$ 1 parties receive a verifiable proof of a party’s proposal, thus ensuring the necessary input for the ABBA instance.
@@
At least one party’s proposal reaches $12f+$ 1 parties.
@@
Since the selected parties can be byzantine and the adversary can schedule the message delivery to delay the agreement, we have considered the following scenarios:
@@
    Among $\langle f+1\rangle$ selected parties, $f$ parties are non-responsive.
@@
    Selected $\langle f+1\rangle$ parties are responsive, but other $f$ non-selected parties are non-responsive.
@@
    Every party is responsive, including the selected $\langle f+1\rangle$ parties.
  4. 4.
    Selected $t<\langle f+1\rangle$ parties are responsive, and total $m$ parties responsive, where $\langle 2f+1\rangle\leq m\leq n$.
We prove that in every scenario, at least one party’s proposal reaches $\langle 2f+1\rangle$ parties.
@@
    Since among $\langle f+1\rangle$ selected parties, $f$ parties are non-responsive, only one party completes the $pVCBC$ protocol and broadcasts the verifiable proof. If any party receives a verifiable proof, then the verifiable proof is from the responsive selected party. Since every party receives the verifiable proof for the same party’s proposal, the proposal reaches $\langle 2f+1\rangle$ parties.
@@
    The $\langle f+1\rangle$ selected parties are responsive and complete the $pVCBC$ protocol. Each selected party broadcasts the verifiable proof, and $\langle 2f+1\rangle$ parties receive the verifiable proof ( another $f$ number of parties are non-responsive). Any party receives a verifiable proof, recommends the received verifiable proof, and waits for $\langle 2f+1\rangle$ recommendations. If a party receives $\langle 2f+1\rangle$ recommendations, then these recommendations include all $\langle f+1\rangle$ parties’ verifiable proofs because among the $\langle 2f+1\rangle$ received recommendations, $\langle f+1\rangle$ number of recommendations are from the selected parties. So, every proposal reaches $\langle 2f+1\rangle$ parties.
@@
    The proof is by contradiction. Let no proposal reaches more than $2f$ parties. Since we assume every party is responsive, every party receives a proposal in $propose$ step. We also know that no party participate in $Sequential-ABBA$ protocol until the party receives $\langle 2f+1\rangle$ recommendations and get the order to Sequential-ABBA. There must be a $\langle 3f+1\rangle*\langle 2f+1\rangle$ recommendation messages. If no proposal can be recommended to more than $2f$ parties, the total number of recommendations is $\langle 3f+1\rangle*\langle 2f\rangle<\langle 2f+1\rangle*\langle 2f+1\rangle$ (Though a proposal can be recommended by more than one party we assume that every party recommends to the same $2f$ parties otherwise it would fulfil the requirement of $\langle 2f+1\rangle$ proposals). But, the honest parties must send enough recommendation messages to ensure the protocol’s progress and the adversary delivers the messages eventually. Therefore, at least one party’s proposal reaches $\langle 2f+1\rangle$ parties, a contradiction.
  4. 4.
    The proof is by contradiction. Let no proposal reaches to more than $2f$ parties. If $1\leq t\leq\langle f+1\rangle$ parties distribute their verifiable proof to $\langle 2f+1\rangle\leq m<\langle 3f+1\rangle$ parties and no proposal reaches more than $2f$ parties, then there must be no more than $m*2f$ recommendations. However, $m$ parties must receive $m$*$\langle 2f+1\rangle$ recommendations greater than $m*2f$, a contradiction.
@@
This lemma ensures that if $12f+$ 1 parties receive a verifiable proof, at least one honest party will input 1 to the ABBA instance, ensuring the protocol reaches an agreement on 1.
@@
The proof works by counting the total number $A$ of $vote=0$ messages (line 44 of Algorithm 4) that are generated by honest parties (over all 00 vote in the agreement loop). Since every honest party has received a verifiable proof from one party in the $propose$ broadcast step (line 18 of Algorithm 3), the party generates $vote=0$ messages for at most $f$ number of selected parties. Thus $A\leq(f)(n-f)$ (all $(n-f)$ votes are zero). For the asynchronous binary byzantine agreement protocol to decide 00 for a particular party’s proposal and to cause one more iteration of the loop, at least $\langle f+1\rangle$ honest parties must propose 00 for the binary agreement (line 46 of Algorithm 4). Since honest parties only propose 00 if they have received $\langle 2f+1\rangle$ $vote=0$ messages, there must be at least $\langle f+1\rangle$ honest parties who have generated a $vote=0$ message in this iteration. Let $R$ denote the number of iterations of the loop where the binary agreement protocol decides 00. From the preceding argument, we have $A\geq R(f+1)$. Combining these two bounds on $A$, we obtain $R(f+1)\leq f*(n-f)$.
$\begin{split}R\leq\frac{(f)(n-f)}{f+1}\\ R\leq\frac{(f)(2f+1)}{f+1}\\ R\leq\frac{(f)(f+1+f)}{f+1}\\ R\leq f+\frac{f^{2}}{f+1}\newline \end{split}$ (1)
However, there are only $\langle f+1\rangle$ number of selected parties’ proposals, and from Lemma 5.1, at least one verifiable proof reaches $\langle f+1\rangle$ honest parties. Therefore, $R<\langle f+1\rangle$ and the adversary can cause at most $R+1$ or $\langle f+1\rangle$ number of iterations. ∎
@@
The binary agreement protocol biased towards $1$ decides 00 if the honest parties input 00. From the Sequential-ABBA protocol, each party must receive $\langle n-f\rangle$ $vote=0$ messages in order to input 00 to an ABBA instance and to decide 00. However, from Lemma 5.1 at least $\langle f+1\rangle$ honest parties receive at least one verifiable proof from a selected party $p$, then it is not the case for the party $p\in\overline{A}$. This proves the first claim.
To prove the second claim, let $A$ denote the total number of 00-votes cast by honest parties. Since an honest party votes 00 for at most $f$ number of parties proposals, we have $A\leq f(n-f)$. On the other hand, a party can cast $vote=0$ for $f+1-|\overline{A}|$ parties proposals, and at least $\langle f+1\rangle$ honest parties cast $vote=0$ to decide 00 then $A\geq(f+1$ $|\overline{A}|)(f+1)$. These bounds on $A$ are the same as in Lemma 5.2 with $R=f+1$-$|\overline{A}|$. Using the same argument, it follows that $|\overline{A}|\geq 1$, since $R$ is at most $f$ [^1] To prove the Lemma 5.2 and Lemma 5.3, we have used the same technique of Lemma 7 and Lemma 9 from [13]. We also refer interested readers to Lemma 10 of the same paper for the pseudorandom generation..
The third claim follows now because $|\overline{A}|$ is at least a constant fraction of $f+1$ and thus, there is a constant $\beta>1$ such that $Pr[\Pi(i))\notin\overline{A}]\leq\frac{1}{\beta}$ for all $1\leq i\leq t$. Since the probability of the $t$ first elements of $\Pi$ jointly satisfying the condition is no larger than for $t$ independently and uniformly chosen values, we obtain $Pr[\Pi[1]\notin\overline{A}\wedge\Pi[2]\notin\overline{A}\wedge....\wedge\Pi[t]\notin\overline{A}]\leq\beta^{-t}$.
@@
Agreement: If an honest party outputs a value $v$, then every honest party outputs the value $v$. From Lemma 5.1, at least one proposal reaches $12f+$ 1 parties, and the ABBA protocol reaches an agreement on $1$ if $f+1$ honest parties input $1$. The agreement property of the ABBA protocol ensures all honest parties output $1$ for that ABBA instance and receive the same value $v$, which satisfies $threshold-validate\langle v,\sigma\rangle=true$.
Liveness: If honest parties participate and deliver messages, all honest parties decide. From Lemma 5.1, at least one proposal reaches $12f+$ 1 parties, ensuring the ABBA protocol decides $1$. Lemma 5.3 confirms the protocol reaches an agreement after a constant expected number of iterations.
External-validity: If an honest party terminates and decides on a value $v$, then $externally-valid\langle v,\sigma\rangle=true$. The validity of the ABBA protocol ensures at least one honest party inputs $1$ to the ABBA instance, meaning the honest party received a valid $threshold-signature$ $\sigma$ for $v$.
Integrity: If honest parties decide on a value $v$, then $v$ was proposed by a party. The ABBA protocol returns $1$ if at least one honest party inputs $1$, which requires a valid $threshold-signature$ $\sigma$. Honest parties reply with $sign-shares$ only if a value $v$ is proposed, ensuring $v$ was proposed by a party. ∎

### v2

@@ 38             $\langle bit,\rho\rangle\leftarrow ABBA\langle v\rangle$ biased towards 11
46              upon receiving $\langle ID,VOTE,p_{k},u_{index^{\prime}},m\rangle$ for the first time from party $p_{k}$ do
47                 if $u_{index^{\prime}}=1$ then
48                   $u_{index}\leftarrow 1$
49                   $m_{index}\leftarrow m$
50
51
52                  $\Sigma\leftarrow\Sigma\cup 1$
53
54
Algorithm 4 Protocol for party $p_{i}$
4.7 Integration of Subprotocols
The pMVBA protocol achieves an agreement in a Byzantine environment through a sequence of interconnected steps. The process begins with the Committee Selection (CS) protocol, where each party dynamically and randomly selects committee members (Algorithm 5, line 10). Following this, the selected parties broadcast their proposals using the Prioritized Verifiable Consistent Broadcast (pVCBC) protocol, ensuring that their broadcasts are verifiable (Algorithm 5, line 13). After the proposals are broadcast, the parties proceed with the Propose-Recommend phase, where verifiable proofs are shared, and recommendations are collected (Algorithm 5, lines 23-27). Upon receiving the necessary recommendations, a random order of the selected committee members is generated using a cryptographic coin-tossing scheme (Algorithm 5, line 18). Finally, the Sequential-ABBA protocol is executed, where the agreement process is run in the determined order until an agreement is reached on a proposal (Algorithm 5, lines 20-21). Each of these steps ensures the protocol operates efficiently and securely, even in the presence of Byzantine faults.
@@ 4.7 Integration of Subprotocols
2Local variables initialization:
@@ 3   $INSTANCE\leftarrow 1$
4   $SelectedParties\leftarrow\bot$
5   $\rho(threshold-signature)\leftarrow\bot$
@@ 6   $Proposal\leftarrow\bot$
7   $RecommendationList\leftarrow\bot$
8   $PermutationList\leftarrow\bot$
@@ 12     $ID\leftarrow{\langle INSTANCE,id\rangle}$
13     $SelectedParties\leftarrow CS\langle ID,INSTANCE\rangle$
@@
15    if $p_{i}\in SelectedParties$ then
16       $Proposal\leftarrow requests$
17       $\rho\leftarrow pVCBC\langle ID,Proposal\rangle$
18       $\textbf{send}\langle PROPOSE,ID,\rho,Proposal\rangle$ to all parties
@@
20       wait until $|RecommendationList|\geq 1$
21
22
@@ 23      Permutation:
24       $PermutationList\leftarrow RandomOrder\langle ID,SelectedParties\rangle$
25
@@ 26      Sequential-ABBA:
27       $Proposal\leftarrow Sequential$-$ABBA\langle PermutationList,RecommendationList\rangle$
@@ 29       $INSTANCE\leftarrow INSTANCE+1$
30
31
32      upon receiving $\langle\textit{PROPOSE},ID_{k},\rho,Proposal\rangle$ for the first time do
33         $RecommendationList\leftarrow SCR\langle ID,ID_{k},Proposal,\rho\rangle$
34
35
36        upon receiving $\langle\textit{RECOMMEND},ID_{k},Proposal,\rho\rangle$ for the first time do
37           if no RECOMMENDATIONRECOMMENDATION is sent yet then
38             $RecommendationList\leftarrow SCR\langle ID,ID_{k},Proposal,\rho\rangle$
@@ 5 Evaluation and Analysis of the Proposed Protocol
• What are the formal proofs supporting the correctness of the pMVBA protocol? ($\S$ 5.1)
• How does the pMVBA protocol perform in terms of message complexity, communication complexity, and running time? ($\S$ 5.2)
• What are the behaviors and outcomes of the protocol under different network conditions and adversarial strategies? ($\S$ 5.3)
• How does the performance of the pMVBA protocol compare to other MVBA protocols in terms of key metrics? ($\S$ 5.4)
• How does pMVBA’s performance compare to atomic broadcast protocols in communication complexity? ($\S$ B.1)
@@ 5.1 Proof of the Prioritized-MVBA Protocol
The correctness of the proposed protocol is critical to ensure that it adheres to the standard MVBA outcomes. Our goal is to prove that reducing the number of broadcasts does not compromise the protocol’s reliability. The output of a party’s request depends on the output of an ABBA instance within our protocol’s agreement loop, typically resulting in a value of 1. For the ABBA protocol to output 1, at least one honest party must input 1. We substantiate our protocol’s integrity by demonstrating that at least $2f+1$ parties receive a verifiable proof of a party’s proposal, thus ensuring the necessary input for the ABBA instance.
@@
At least one party’s proposal reaches $2f+1$ parties.
@@
Since both the selected and non-selected parties can be non-responsive (Byzantine or system failure), and the adversary can coordinate with the Byzantine parties to schedule the message delivery in a way that slows down the agreement process. To prove the lemma, we parameterized the number of responsive parties among the selected parties and the total number of responsive parties among both selected and non-selected parties. Below is the parameterized statement.
  • Among the $f+1$ selected parties, $t$ parties are responsive where $1\leq t\leq f+1$ , and among total $n$ parties, $m$ number of parties are responsive, where $2f+1\leq m\leq n$. We provide three scenarios below, which are special cases of the above statement. We utilize the special cases to prove the parameterized statement. The special cases are three edge cases where the maximum number of selected and the non-selected parties can be responsive and non-responsive (case 1 and 2)
@@
      Among $f+1$ selected parties, $f$ parties are non-responsive. (The maximum number of selected parties can be non-responsive)
@@
      Selected $f+1$ parties are responsive, but other $f$ non-selected parties are non-responsive. (The maximum number of non-selected parties can be non-responsive)
@@
      Every party is responsive, including the selected $f+1$ parties. (The maximum number of responsive parties)
    We will first prove that the above three scenarios are a special case of the parameterized statement. These three cases are the lower and upper bound for $t$ and $m$.
@@
      For case 1, $t=1$ (selected responsive parties) and $m=t+2f=2f+1$ (since $f$ selected parties are non-responsive, the total number of responsive parties is $2f+1$ that includes one responsive selected party). So it adheres requirement of the parameterized equation.
@@
      For case 2, $t=f+1$ and m=t+f=f+1+f(f+1selectedresponsivepartyandfnonseletedresponsiveparty)=2f+1m=t+f=f+1+f(f+1selectedresponsivepartyandfnon-seletedresponsiveparty)=2f+1. So it also adheres requirement of the parameterized equation.
@@
      For case 3, $t=f+1$ and $m=t+2f=3f+1=n$. Therefore, it also adheres requirement of the parameterized equation.
We prove that in every scenario, at least one party’s proposal reaches $\langle 2f+1\rangle$ parties. Since we have proved that cases (1), (2), and (3) are the special cases of a parameterized statement, we utilize these cases to prove that even in the scenario of parameterized statement, at least one proposal reaches $2f+1$ parties.
  1. 1.
    The proof proceeds by contradiction. Assume that among the $f+1$ selected parties, $f$ number of parties are nonresponsive, meaning only one party completes the $pVCBC$ protocol and broadcasts the verifiable proof. Under the assumption that no proposal reaches more than $2f$ parties, the total number of recommendations would be:
    $(1)(2f)(\frac{2f+1}{1})$
    On average, a selected party’s proposal is recommended by $\frac{2f+1}{1}$ parties. Simplifying this:
[…]
    Thus, the total number of recommendations is $(2f)(2f+1)$, which is less than the required $(2f+1)(2f+1)$ recommendations.
    This leads to a contradiction, as the protocol necessitates a sufficient number of recommendation messages to guarantee progress. Given that honest parties are obligated to provide the necessary number of recommendations, it follows that the proposal must have been received by a minimum of $2f+1$ parties. Consequently, the assumption that no proposal reaches more than $2f$ parties is incorrect, and at least one selected party’s proposal reaches $2f+1$ parties, thereby enabling the protocol to function as intended.
  2. 2.
    We will prove by contradiction. Assume that no proposal reaches more than $2f$ parties.
    Given the assumption that a minimum of $2f+1$ parties are responsive and that $f+1$ of these are chosen to propose, it follows that every party must receive a proposal during the proposepropose step. Furthermore, no party commences the Sequential-ABBA protocol until it has received a minimum of $2f+1$ recommendations and has secured the random order necessary to advance with Sequential-ABBA. This indicates that a total of $(2f+1)(2f+1)$ recommendation messages must be exchanged.
    Assuming that no proposal is recommended to more than $2f$ parties and that $f+1$ parties submit their requests, the total number of recommendations would be:
    $(f+1)(2f)(\frac{2f+1}{f+1})$
    On average, a proposal from a selected party would receive recommendations from $\frac{2f+1}{f+1}$ parties. Simplifying the expression:
[…]
    Therefore, the total number of recommendations is $(2f)(2f+1)$, which is lesser to the necessary $(2f+1)(2f+1)$ recommendation messages.
    This results in a contradiction, as the protocol necessitates the transmission of sufficient recommendation messages for progress to occur. Honest parties are required to ensure the delivery of the necessary number of recommendations; thus, at least one proposal must be received by $2f+1$ parties. Therefore, our assumption that no proposal reaches more than $2f$ parties is incorrect. Consequently, at least one party’s proposal successfully reaches $2f+1$ parties, thereby ensuring the protocol can advance as necessary.
  3. 3.
    The proof proceeds by contradiction. Assume that no proposal reaches more than $2f$ parties.
    Since we assume that all parties are responsive, each party receives a proposal during the proposepropose step. Additionally, no party engages in the Sequential-ABBA protocol until it has received a minimum of $2f+1$ recommendations and acquired the random order required to execute the Sequential-ABBA protocol. Thus, the protocol requires a total of $(3f+1)(2f+1)$ recommendation messages.
    If no proposal is recommended to more than $2f$ parties and $f+1$ parties propose their requests, then the total number of recommendations must not exceed $(f+1)(2f)$. On average, a proposal from a selected party can receive recommendations from $\frac{3f+1}{f+1}$ parties. Therefore, the total number of recommendations does not exceed:
    $\begin{split}(f+1)(2f)(\frac{3f+1}{f+1})\end{split}$ (3)
    Simplifying:
[…]
    Therefore, the total number of recommendations can be expressed as $(2f)(3f+1)$. The necessary quantity of recommendations is $(3f+1)(2f+1)$.
    While a proposal can be recommended by more than one party, if it is assumed that all recommending parties recommend the same set of $2f$ parties, the condition of receiving $2f+1$ recommendations remains unfulfilled. Consequently, for the protocol to advance, it is essential that the honest parties broadcast a sufficient number of recommendation messages, which the adversary will ultimately deliver.
    Therefore, a minimum of one proposal must be communicated to $2f+1$ parties. This contradicts the initial assumption; therefore, at least one party’s proposal has to reach $2f+1$ parties, thereby ensuring the protocol’s advancement.
We prove that in every scenario, at least one party’s proposal has been shared to $\langle 2f+1\rangle$ parties. Having shown that cases (1), (2), and (3) are particular instances of the general parameterized scenario, the proof of these specific cases allow us to extend the result to the parameterized scenario.
In the three scenarios analyzed, it is evident that with $m$ responsive parties, the total number of recommendations are $m(2f+1)$. If $t$ selected parties are responsive, the total number of recommendations can be represented as $t(2f)(\frac{m}{t})$. In this context, $(t)(2f)$ indicates that the proposals from each of the $t$ selected parties can reach a maximum of $2f$ parties. Meanwhile, $\frac{m}{t}$ denotes the average number of parties that each selected party’s proposal can reach.
Therefore, we establish the subsequent relationship:
[…]
Simplifying further:
[…]
This inequality confirms that the number of recommendations from the selected parties is less than the required total number of recommendations, proving that at least one party’s proposal must reach $2f+1$ parties, ensuring protocol progress.
@@
This lemma ensures that if $2f+1$ number of parties receive verifiable proof, at least one honest party will input 1 to the ABBA instance, ensuring the protocol reaches an agreement on 1.
@@
Since only $f+1$ parties are selected, the number of iterations is bounded above by $f+1$. By Lemma 5.1, at least one verifiable proof is delivered to $f+1$ honest parties, which guarantees that the protocol reaches agreement on at least one party’s proposal. We therefore analyze the worst case in which the protocol requires exactly $f+1$ iterations. If the ABBA execution order of the selected parties is known in advance, an adaptive adversary can schedule message deliveries so that only the proposal of the last party in that order attains recommendations from $2f+1$ number of parties. Consequently, for all earlier selected parties, less than the threshold number of parties cast $vote=1$ and the instance decides 0; for the final selected party, at least $f+1$ honest parties cast $vote=1$ and the instance decides $1$. . ∎
@@
The binary agreement protocol biased towards $1$ decides 0 if the honest parties input 0. From the Sequential-ABBA protocol, each party must receive $\langle n-f\rangle$ $vote=0$ messages in order to input 0 to an ABBA instance and to decide 0. However, from Lemma 5.1 at least $\langle f+1\rangle$ honest parties receive at least one verifiable proof from a selected party $p$, then it is not the case for the party $p\in\overline{A}$. This proves the first claim.
To prove the second claim, we refer to the two special cases of Lemma 5.1. In case $1$, one selected party is responsive and every party receives the verifiable proof of that selected party; therefore, $|\overline{A}|=1$. For case 2, every party receives a recommendation for each selected party. Therefore, each selected party’s verifiable proof reaches $2f+1$ parties, thus $|\overline{A}|=f+1$. Since Lemma 5.1 proves that at least one party’s proposal reaches $2f+1$ parties, it generalizes the claim $|\overline{A}|\geq 1$.
The third claim follows now because $|\overline{A}|$ is at least a constant fraction of $f+1$ and thus, there is a constant $\beta>1$ such that $Pr[\Pi(i))\notin\overline{A}]\leq\frac{1}{\beta}$ for all $1\leq i\leq t$. Since the probability of the $t$ first elements of $\Pi$ jointly satisfying the condition is no larger than for $t$ independently and uniformly chosen values, we obtain $Pr[\Pi[1]\notin\overline{A}\wedge\Pi[2]\notin\overline{A}\wedge....\wedge\Pi[t]\notin\overline{A}]\leq\beta^{-t}$  [^1] To prove the part of the Lemma 5.3, we have used the same technique of Lemma 9 from [SECURE03]. We also refer interested readers to Lemma 10 of the same paper for the pseudorandom generation..
@@
Agreement: If an honest party outputs a value $v$, then every honest party outputs the value $v$. From Lemma 5.1, at least one proposal reaches $2f+1$ parties, and the ABBA protocol reaches an agreement on $1$ if $f+1$ honest parties input $1$. The agreement property of the ABBA protocol ensures all honest parties output $1$ for that ABBA instance and receive the same value $v$, which satisfies $threshold-validate\langle v,\rho\rangle=true$.
Liveness: If honest parties participate and deliver messages, all honest parties decide. From Lemma 5.1, at least one proposal reaches $2f+1$ parties, ensuring the ABBA protocol decides $1$. Lemma 5.3 confirms the protocol reaches an agreement after a constant expected number of iterations.
External-validity: If an honest party terminates and decides on a value $v$, then $externally-valid\langle v,\rho\rangle=true$. The validity of the ABBA protocol ensures at least one honest party inputs $1$ to the ABBA instance, meaning the honest party received a valid thresholdsignaturethreshold-signature $\rho$ for $v$.
Integrity: If honest parties decide on a value $v$, then $v$ was proposed by a party. The ABBA protocol returns $1$ if at least one honest party inputs $1$, which requires a valid thresholdsignaturethreshold-signature $\rho$. Honest parties reply with signsharessign-shares only if a value $v$ is proposed, ensuring $v$ was proposed by a party. ∎